# The Cone Restriction: An old Approach Revisited

### v5

@@ 3.2 The second algorithm
Verify the properties Property 2 follows from (3.25) and the repeated application of (3.20). Property 3 follows from (3.26) and the repeated application of (3.21). Property 1 follows from (3.19), (3.37), (3.24), (3.35), and Lemma 2.19. See [11] for details. Here we verify Property 4. Apply (3.23) repeatedly, one has
${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox\prod^{n-2}_{i=l}{D^{\delta}_{i}{\left(\frac{r_{i+1}}{r_{i}}\right)}^{-\frac{n-i-1}{2}}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}=r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}D^{\delta}_{j-1}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$
This gives Property 4.
3.3 Establish the $k$-broad estimates
@@
The task now is how to apply the properties guaranteed by the algorithms [alg 2]. Since each $O\in\mathcal{O}$ has diameter at most $R^{O\left({\delta}_{0}\right)}$, trivially one may bound
${\left\|Ef_{O}\right\|}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}\lessapprox{\left\|f_{O}\right\|}_{2}$ (3.38)
Combine (3.29) and (3.38),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p_{m}}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}}\right)}^{{\beta}_{m}/p_{m}\ }$
@@
$\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\right)}^{{\beta}_{m}/p_{m}\ }{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2{\beta}_{m}\left(\frac{1}{2}-\frac{1}{p_{m}}\right)}_{2}\ }$ (3.39)
Combine (3.30), (3.27), and (3.29),
@@
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{i}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{1-\frac{2}{p}}_{2}\ }$ (3.40)
By taking $l=n$ in (3.32),
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{i=m-1}{r^{-1/2}_{i}D^{-i+\delta}_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.41)
Combine (3.40) and (3.41),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{i}}{\left\|f\right\|}^{2}_{2}$ (3.42)
By choosing suitable parameters $p_{i}$ for $k\leq l\leq n$, one can guarantee
$\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \ \ \ \ \ \ \frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \forall m\leq i\leq n-1$
This gives Theorem 1.3. See [11] for details.
4 Proof of Theorem 2.23
In this section we adapt the method provided in [12] to prove Theorem 2.23. Let $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ and $A$ is a large integer. Without loss of generality, we assume that $f$ satisfies (3.27). Applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described in Section 3.2.
First, we improve (3.42) by Corollary 2.17, the nested polynomial Wolff axioms.
@@
For $m\leq l<n-1$,
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox D^{\delta}_{n-1}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right){\left\|f\right\|}^{2}_{\infty}$
@@
Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right$], Corollary 2.17 implies
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.1)
@@
Put (3.22) into this bound,
@@
${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$ (4.2)
Now Lemma (4.1) follows from (4.1) and (4.2). ∎
Combining (3.32), (3.33), and Lemma 4.1, we see that for $m\leq l<n-1$,
@@
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-\frac{1}{2}}_{i}D^{-i+\delta}_{i}}\right)r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-\frac{1}{2}}_{j}D^{\delta}_{j-1}}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right)D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$
$\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
Let ${\gamma}_{m},{\gamma}_{m+1},\dots,{\gamma}_{n}$ be positive real numbers satisfying ${\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n}=1$, and
${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}},\ \ {\tau}_{j}:=\left\{\begin{array}[]{c}{\sigma}_{j}\mathrm{,\ \ if\ }j\leq n-2\\ 0,\ \ \mathrm{if\ }j=n-1\end{array}\right$.
Take the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$,
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m}{r^{-\frac{1+{\tau}_{j}}{2}}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)+\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$
Put this bound into (3.40),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{j=m}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\tau}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}\prod^{n-1}_{j=m-1}{D^{\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{j}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\left\|f\right\|}^{1-\frac{2}{p}}_{\infty}$
It just remains to find suitable parameters $p_{i}$ for $m\leq i\leq n$ such that
$X_{j}:=\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\tau}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \mathrm{for\ }m\leq j\leq n-1$
$Y_{j}:=\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \mathrm{for\ }m-1\leq j\leq n-1$
To find the critical values, we assume that $X_{j}=0$ for $m\leq j\leq n-1$ and $Y_{j}=0$ for $m-1\leq j\leq n-1$. By (3.27),
$\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j+1}}\right)}^{-1}-\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j}}\right)}^{-1}=\frac{1+{\tau}_{j}}{2},\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.4)
$\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j+1}}\right)}^{-1}=1+j\left(1-{\sigma}_{j}\right),\ \ \mathrm{for\ }m-1\leq j\leq n-1$ (4.5)
Putting (4.5) into (4.4),
$j\left(1-{\sigma}_{j}\right)-\left(j-1\right)\left(1-{\sigma}_{j-1}\right)=\frac{1+{\tau}_{j}}{2},\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.6)
Thus, ${\gamma}_{m},{\gamma}_{m+1}\dots,{\gamma}_{n-1}$ can be determined recursively. Since the initial step and the recursive steps for $j\leq n-2$ are same as that in [12], we have
@@
${\gamma}_{j}=\frac{m-1}{2}\frac{1}{j\left(j-1\right)}\prod^{j}_{i=m}{\frac{2i}{2i+1}},\ \ \mathrm{for\ }m\leq j\leq n-2$ (4.7)
Since the relationship (4.5) and ${\gamma}_{m},{\gamma}_{m+1}\dots,{\gamma}_{n-2}$ are same as that in [12], one has
@@
$p_{j}=2+\frac{6}{2\left(j-1\right)+\left(m-1\right)\prod^{j-1}_{i=m}{\frac{2i}{2i+1}}}\ ,\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.8)
The condition $X_{n-1}=0$ is equivalent to
${\left(\frac{1}{2}-\frac{1}{p_{n}}\right)}^{-1}-{\left(\frac{1}{2}-\frac{1}{p_{n-1}}\right)}^{-1}=1$ (4.9)
By putting (4.8) into (4.9), one has
$p_{n}=2+\frac{12}{4n-5+2\left(m-1\right)\prod^{n-2}_{i=m}{\frac{2i}{2i+1}}}$
By remark 3.2, $m\geq k$. Thus, if we take
$p>2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$
Then the broad estimate always (2.7) holds. It just remains to check ${\sigma}_{n-1}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-1}\leq 1$. By (4.5),
@@
${\sigma}_{n-2}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-2}<\frac{m-1}{2}\sum_{j\geq m}{\frac{1}{j\left(j-1\right)}}=\frac{1}{2}$ (4.10)
Take j=n1j=n-1 in (4.6),
$\left(n-1\right)\left(1-{\sigma}_{n-1}\right)-\left(n-2\right)\left(1-{\sigma}_{n-2}\right)=1/2\ \Longrightarrow{\gamma}_{n-1}=\frac{1/2-{\sigma}_{n-2}}{n-1}$
By (4.10),
${\gamma}_{n-1}<\frac{1}{2\left(n-1\right)}$ (4.11)
Combine (4.10) and (4.11), we see that ${\sigma}_{n-1}<1$.

### v7

@@ 3.2. Algorithm 2
Verification of the Properties. Property 2 follows from (3.22) together with repeated applications of (3.17). Property 3 follows from (3.23) and repeated applications of (3.18). Property 1 follows from (3.16), (3.34), (3.21), (3.33), Lemmas 2.12 and 2.13; see [11] for details. We now verify Property 4.
Applying (3.20) repeatedly yields
$\|f_{\overrightarrow{S}_{l}}\|_{2}^{2}\lessapprox\prod_{i=l}^{n-2}\left(\frac{r_{i+1}}{r_{i}}\right)^{-(n-i-1)/2}\|f^{\#}_{\overrightarrow{S}_{l}}\|_{2}^{2}=r_{l}^{(n-l-1)/2}\prod_{j=l+1}^{n-1}r_{j}^{-1/2}\|f^{\#}_{\overrightarrow{S}_{l}}\|_{2}^{2}.$
This establishes Property 4.
3.3. Establish the Broad Estimates
@@
We now apply the properties provided by Algorithm 2. Since each $O\in\mathcal{O}$ has diameter at most $R^{O(\delta_{0})}$, we may trivially bound
${\left\|Ef_{O}\right\|}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}\lessapprox{\left\|f_{O}\right\|}_{2}$
Together with (3.27), this yields
@@
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\right)}^{{\beta}_{m}/p_{m}\ }{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2{\beta}_{m}\left(\frac{1}{2}-\frac{1}{p_{m}}\right)}_{2}\ }$ (3.35)
Combining this with (3.28) and (3.25), we obtain
@@
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(\frac{1}{2}-\frac{1}{p}\right)}_{i}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{1-\frac{2}{p}}_{2}\ }$ (3.36)
4. Complete the proof
In this section, we complete the proof of Proposition 2.16. For the reader’s convenience, we restate the proposition below.
Proposition 4.1.
Let $n\geq 3$ and $k\in[2,n-2]$. For
$p=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}},$
$\frac{1}{q}=\frac{1}{p}+\frac{1}{4n-8}\big(\frac{6}{p}-1\big),$
and some integer $A\lesssim_{\varepsilon,n}1$, the estimate
$\|Ef\|_{BL_{k,A}^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$
holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
Suppose that, after applying Algorithm 2 to decompose the $k$-broad norm $\|Ef\|_{BL^{p}_{k,A}(B_{R})}$, we obtain the outputs described in Section 3.2. By Lemma 2.14, we have $m\geq k$. Consequently, we divide the proof into cases $n-1\leq m\leq n$ and $k\leq m\leq n-2$.
Case $n-1\leq m\leq n$ . Taking $l=n$ in (3.30), we obtain
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m-1}{r^{-1/2}_{j}D^{-j}_{j}}\right){\left\|f\right\|}^{2}_{2}$ (4.1)
Combining this with (3.36), we have
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{j=m-1}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}D^{\frac{{\beta}_{j+1}}{2}-\left(j+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}{\left\|f\right\|}^{2}_{2}$ (4.2)
Set
$\begin{cases}p=p_{n}=\frac{2n}{n-1}&\text{if $m=n$},\\[4.0pt] p_{n-1}=\frac{2(n-1)}{n-2},\quad p=p_{n}=\frac{2(2n-1)}{2n-3}&\text{if $m=n-1$}.\end{cases}$ (4.3)
In both cases, $p<p(n,k)$ and
$\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0,\quad\forall j\in[m-1,n-1]$
$\frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\quad\forall j\in[m-1,n-1]$
This finishes the proof in this case.
Case $k\leq m\leq n-2$ . Applying the nested polynomial Wolff axioms, we have
@@ 4. Complete the proof
For $m\leq l\leq n-2$,
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right){\left\|f\right\|}^{2}_{\infty}$
@@
Since $f^{\#}_{\overrightarrow{S}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $(\theta,v,l)$ with $\theta\in\Theta[\overrightarrow{S}^{\#}_{l}$], Corollary 2.27 implies
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\\ \lessapprox\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.4)
@@
Together with (3.19) we have
@@
${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.5)
Lemma (4.2) now follows from (4.4) and (4.5). ∎
Combining (3.30), (3.31), and Lemma 4.2, we obtain that for $m\leq l\leq n-2$,
@@
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.6)
Define $p_{j}$ by
$p_{j}:=\begin{cases}2+\frac{6}{2\left(j-1\right)+\left(m-1\right)\prod^{j-1}_{i=m}{\frac{2i}{2i+1}}}&\text{if $m\leq j\leq n-1$},\\[4.0pt] p(n,k)&\text{if $j=n$}.\end{cases}$ (4.7)
For $m\leq j\leq n-1$, we define $\gamma_{j}\in[0,1]$ by
@@
$\gamma_{j}:=\begin{cases}\frac{m-1}{2}\frac{1}{j\left(j-1\right)}\prod^{j}_{i=m}{\frac{2i}{2i+1}}&\text{if $m\leq j\leq n-2$},\\[4.0pt] 1-\sigma_{n-2}&\text{if $j=n-1$}.\end{cases}$ (4.8)
where
${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}}$
Taking the geometric mean of (4.6) with weights $\gamma_{l}$ and (4.1) with weight $\gamma_{n-1}$, we obtain
@@
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\\ {r_{n-1}^{-\frac{1}{2}}}D_{n-1}^{-\gamma_{n-1}(n-1)}\left(\prod^{n-2}_{j=m}{r^{-\frac{1+{\sigma}_{j}}{2}}_{j}}\right)\left(\prod^{n-2}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)}_{j}}\right){\left\|f\right\|}^{2\gamma_{n-1}}_{2}{\left\|f\right\|}^{2\sigma_{n-2}}_{\infty}$
Together with (3.36), we see that
@@
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox{r^{\frac{{\beta}_{n}-{\beta}_{n-1}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{n-1}}D^{\frac{{\beta}_{n}}{2}-\left(1+\gamma_{n-1}(n-1)\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{n-1}\\ \prod^{n-2}_{j=m}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\sigma}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}\prod^{n-2}_{j=m-1}{D^{\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}{\left\|f\right\|}^{\frac{2}{q}}_{2}{\left\|f\right\|}^{1-\frac{2}{q}}_{\infty}$
where $q$ is defined by
$\frac{1}{q}=\frac{1}{p}+\frac{1}{4n-8}\big(\frac{6}{p}-1\big)$
Arguing as in [12], we see that
$\frac{{\beta}_{n}-{\beta}_{n-1}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0$
$\frac{{\beta}_{n}}{2}-\left(1+\gamma_{n-1}(n-1)\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0$
$\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\sigma}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0,\qquad\forall j\in[m,n-2]$
$\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)=0,\qquad\forall j\in[m-1,n-2]$
This concludes the proof since $k\leq m$.